# Climate impacts of equatorial superrotation

### v1

@@ 1 Introduction
In the context of anthropogenic climate change, the response of the atmospheric circulation to radiative forcing is often seen as a second-order effect, even if it is widely acknowledged that it can have large consequences localized in time and space, for instance for regional climate and extreme events (Cohen et al., 2014; Shepherd, 2014; Vautard et al., 2023, e.g.). In principle, it is possible that in a warmer climate a major reorganization of the large-scale circulation would drive further global changes, either through positive feedbacks acting uniformly in space or by changing the energy balance of large portions of the globe.
@@
While fundamental balance relations enforce strong constraints on the large-scale mid-latitude circulation, dramatic changes are not a priori excluded for the tropical circulation, and it has been suggested that major surprises might occur there, potentially through coupling with the ocean (Pierrehumbert, 2000). While current projections do exhibit changes for instance in the extent of the Hadley cell (Seidel et al., 2008), the intertropical convergence zone (Byrne et al., 2018), and a general weakening of the overturning circulation (Merlis, 2015; Chemke and Yuval, 2023), they do not foresee major changes in the structure of the circulation. On the other hand, the emergence of equatorial superrotation – the reversal of Earth’s climatological easterly winds in the tropics (Held, 1999) – would constitute a fundamental change in the atmospheric general circulation. Superrotation has long been studied by atmospheric dynamicists and modelers (Suarez and Duffy, 1992; Saravanan, 1993; Held, 1999), but has attracted renewed interest in recent years partly because of its relevance for planetary atmospheres (Read and Lebonnois, 2018; Imamura et al., 2020). The suggestion that equatorial superrotation might have occurred during warm climates of Earth’s past, such as those of the Pliocene (Tziperman and Farrell, 2009), Early Eocene (Caballero and Huber, 2010), or even deeper in time in the Mesozoic (Lan et al., 2023), has further motivated study of this atmospheric phenomenon.
@@
Until now, the focus of superrotation research has been on the dynamical processes sustaining equatorial westerlies, as this requires non-trivial anti-diffusive angular momentum transport mechanisms (Held, 1999). Several routes have been suggested, all involving momentum transport by equatorial waves, either through enhanced diabatic heating (Kraucunas and Hartmann, 2005; Arnold et al., 2012; Lutsko, 2018) or through a hydrodynamic instability (Williams, 2003; Mitchell and Vallis, 2010; Potter et al., 2014; Zurita-Gotor and Held, 2018; Zurita-Gotor et al., 2022; Zurita-Gotor and Held, 2024). It has further been suggested that those mechanisms might include positive feedbacks leading to an abrupt transition (Arnold et al., 2012; Herbert et al., 2020; Zurita-Gotor et al., 2022), which might make superrotation an example of a tipping point. Most of these dynamical studies were conducted in an idealized modeling framework of a dry atmosphere without explicitly taking into account processes including condensation of water vapor and radiation. Two exceptions are the studies of Caballero and Huber (2010), which observed a spontaneous transition to superrotation under large radiative forcing, and Arnold et al. (2013), which reported weak superrotation in response to increasing sea surface temperature (SST).
@@
In this work, rather than focusing on the dynamical processes underlying superrotation, we instead explore the consequences that such a large-scale circulation change would have for climate if it were to occur, in particular for surface temperature and the water cycle. To do so, we impose superrotation using a prescribed torque in a moist model which includes parameterizations for major physical processes in the atmosphere such as radiation, convection, and condensation (see Section 2 for details). We find that superrotation is not restricted to a zonal wind change in the upper tropical troposphere but rather affects the general circulation as a whole, resulting in substantial impacts on surface temperature and precipitation. These impacts could potentially be important for interpreting paleoclimate proxies and the processes controlling warm climates in Earth’s past and future. The impact of (forced) superrotation on surface temperature, revealed in this study, together with the emergence of superrotation in simulations of hot climates  (Caballero and Huber, 2010; Arnold et al., 2013) suggests a possible feedback between superrotation and SST that has received little attention to date.
@@
After presenting the modeling setup used in our study (Section  2), we discuss the temperature and precipitation changes induced by superrotation (Section  3). We then build an understanding of the physical processes underlying the temperature changes by investigating the top-of-atmosphere (TOA) radiative balance using a forcing/feedback framework (Section  4) and reconstructing the simulated temperature response using an energy balance model (Section  5). Finally, we relate the superrotation-induced water cycle changes to the circulation changes (Section  6) before summarizing our results (Section  7).
@@ 2.1 Numerical experiments
In this study, we use the GCM Isca (Vallis et al., 2018) in a relatively idealized moist aquaplanet configuration. The model, based on the GFDL dynamical core (Gordon and Stern, 1982), solves the moist primitive equations in vorticity-divergence form on 40 (unevenly spaced) $\sigma$ levels (Bourke, 1974) at T42 horizontal resolution.
@@
Latent heat of condensation and the radiative effects of water vapor are simulated, with the former represented through the large-scale condensation parameterization of Frierson et al. (2006). A standard Betts-Miller scheme for moist convection is also employed (Betts, 1986; Betts and Miller, 1986). Radiative fluxes are computed using the comprehensive SOCRATES scheme (Manners et al., 2017), with a CO2222 concentration of $300$ ppm and a modern ozone climatology. There is a diurnal cycle in the model but no seasonal cycle; a permanent equinox configuration is used for insolation. Clouds are not represented in our configuration of the model and the radiative effect of aerosols is not taken into account.
@@
The lower boundary is a global slab ocean with a $20$ m mixed layer depth. We impose prescribed meridional heat fluxes in the ocean, following Merlis et al. (2013). Sea ice is not represented in the model and the surface temperature is allowed to reach values below $0^{\circ}$. The surface albedo is prescribed to $0.25$ and standard drag laws govern the exchanges of momentum, sensible and latent heat at the surface, based on Monin-Obukhov theory as described in Frierson et al. (2006).
@@
This type of idealized configuration (with variations regarding, for example, how radiative transfer is treated or whether prescribed or interactive SSTs are used) has been used extensively, not only as an evaluation framework for model intercomparison (Neale and Hoskins, 2000a, b; Lee et al., 2008) but also as a conceptual framework to connect theory with numerical models which remain interpretable (Blackburn and Hoskins, 2013). For instance, similar configurations have been used to study meridional energy transport (Caballero and Langen, 2005; Frierson et al., 2007; Rose and Ferreira, 2012; Merlis et al., 2022) and the water cycle (O’Gorman and Schneider, 2008; Byrne and Schneider, 2016; Donohoe et al., 2019) over a range of climates, the response of the Hadley cells to various forcings (Levine and Schneider, 2011; Merlis et al., 2013; Hilgenbrink and Hartmann, 2018), convectively coupled waves (Frierson, 2007), and the annual cycle of the tropical tropopause layer (Jucker and Gerber, 2017).
Here, we use the idealized configuration described above because it constitutes a relatively simple framework that nevertheless takes into account key feedbacks linking surface temperature and the water cycle to changes in large-scale circulation. The strategy we adopt is to impose a superrotating state by prescribing in the upper troposphere eddy zonal momentum fluxes with the following form:
$\mathbf{F}=F_{0}\exp\left(-\frac{1}{2}\left(\frac{\phi}{\phi_{\rm scale}}\right)^{2}\right)\exp\left(-\frac{1}{2}\left(\frac{p-p_{0}}{p_{\rm scale}}\right)^{2}\right)\mathbf{e}_{x},$ (1)
where $\phi_{\rm scale}=5^{\circ}$, $p_{0}=300$ hPa, $p_{\rm scale}=100$ hPa and $\mathbf{e}_{x}$ is the zonal unit vector. We vary the strength of the prescribed torque through the $F_{0}$ parameter, from 00 to $4$ m.s11-1- $-1$, so as to force a transition into a state of equatorial superrotation. The precise mechanisms generating such eddy fluxes are not the focus of this study, but previous work has illustrated cases where eddy momentum flux convergence was generated by zonally asymmetric diabatic heating (Kraucunas and Hartmann, 2005; Arnold et al., 2012; Lutsko, 2018), by a low-latitude mixed barotropic-baroclinic instability (Williams, 2003), or by a Kelvin-Rossby instability (Mitchell and Vallis, 2010; Potter et al., 2014; Zurita-Gotor and Held, 2018; Zurita-Gotor et al., 2022; Zurita-Gotor and Held, 2024).
@@ 2.2 Forcing/feedback framework
We run several experiments with the model described above and interactive SSTs: a control run, with $F_{0}=0$ m.s11-1- $-1$, as well as forced runs with unit increments of $F_{0}$ up to $F_{0}=4$ m.s11-1- $-1$ (we will mostly focus on the $F_{0}=0$ and $F_{0}=4$ m.s11-1- $-1$ experiments). Analogous runs but with fixed SSTs are also performed: in those cases, we prescribe time-averaged SSTs from the control run. The fixed-SST experiments are run in order to diagnose the radiative forcing induced by the prescribed superrotation, in the standard way described for instance by Dessler and Zelinka (2015). In particular, our aim is to decompose the changes in TOA radiation into forcing and feedback components, which can each be broken down into contributions from different physical mechanisms. In this context, we can write the anomalous TOA energy budget as:
$\underbrace{\delta R}_{\text{Total TOA balance}}=\underbrace{F}_{\text{Forcing}}-\underbrace{\lambda\delta T_{\rm s}}_{\text{Feedbacks}},$ (2)
@@
The radiative forcing term $F$ in Eq. (2), diagnosed using the fixed-SST simulations, can be interpreted as the change in the TOA radiation budget driven by the prescribed torque and circulation change. The feedback term $\lambda\delta T_{\rm s}$, on the other hand, includes all changes in TOA radiation linearly related to surface temperature. This is the standard framework for studying the climate sensitivity to CO2 radiative forcing (Sherwood et al., 2020). However, it is also applicable to our study where the question is not the climate response to CO2 forcing but rather the climate response to an imposed superrotating state, which we expect to drive a radiative forcing through a reorganization of water vapor and temperature. In particular, we are interested in distinguishing the direct consequences of the circulation changes we impose, through the prescribed torque at the equator, from the consequences related to feedbacks involving water vapor and temperature.
@@ 2.3 Radiative kernels
We decompose changes in the TOA radiation budget into components representing different physical processes using radiative kernels (Soden et al., 2008). Kernels quantify the radiative responses of the system to a unit change of a given quantity at a given point in space. In general terms and assuming linearity, we can approximately write a TOA radiation change, $\delta R$, as a function of the radiative kernels and changes in the various physical quantities affecting TOA radiation:
$\delta R\approx\sum_{X}\underbrace{\frac{\partial R}{\partial X}}_{\text{radiative kernel}}\delta X,$ (3)
where $\partial R/\partial X$ is a radiative kernel and $X$ represents quantities affecting the TOA radiation, such as temperature or specific humidity. Kernels are computed by running a radiation scheme on two base states that differ only by a local change in the chosen variable, for instance a $1$ K temperature change or a $1$ specific humidity change at a given point in space. In our case, we use radiative kernels computed for Isca with the same setup as that employed in this study (aquaplanet, no clouds or sea ice, T42424242 resolution, $40$ vertical levels) (Liu, 2020), but with the RRTM radiative scheme (Mlawer et al., 1997) (rather than SOCRATES). To validate this approach, we have verified that the total radiative budget computed from the kernels approximately matches the simulated radiative budget from the experiments (see Section 4.4.2).

### v3

# Climate impacts of forced equatorial superrotation in an idealized GCM

@@ 1 Introduction
In the context of anthropogenic climate change, the response of the atmospheric circulation to radiative forcing is often seen as a second-order effect, even if it is widely acknowledged that it can have large consequences localized in time and space, for instance for regional climate and extreme events (Cohen2014; Shepherd2014; Vautard2023, e.g.). In principle, it is possible that in a warmer climate a major reorganization of the large-scale circulation would drive further global changes, either through positive feedbacks acting uniformly in space or by changing the energy budget of large portions of the globe.
@@
While fundamental balance relations enforce strong constraints on the large-scale mid-latitude circulation, dramatic changes are not a priori excluded for the tropical circulation, and it has been suggested that major surprises might occur there, potentially through coupling with the ocean (pierrehumbertClimateChangeTropical2000). While current projections do exhibit changes for instance in the extent of the Hadley cell (Seidel2008), the intertropical convergence zone (Byrne2018), and a general weakening of the overturning circulation (Merlis2015; Chemke2023), they do not foresee major changes in the structure of the circulation. On the other hand, the emergence of equatorial superrotation – the reversal of Earth’s climatological easterly winds in the tropics (heldEquatorialSuperrotationEarthlike1999) – would constitute a fundamental change in the atmospheric general circulation. Superrotation has long been studied by atmospheric dynamicists and modelers (suarezTerrestrialSuperrotationBifurcation1992; saravananEquatorialSuperrotationMaintenance1993; heldEquatorialSuperrotationEarthlike1999), but has attracted renewed interest in recent years partly because of its relevance for planetary atmospheres (readSuperrotationVenusTitan2018; Imamura2020). The suggestion that equatorial superrotation might have occurred during warm climates of Earth’s past, such as those of the Pliocene (tzipermanPlioceneEquatorialTemperature2009b), Early Eocene (caballeroSpontaneousTransitionSuperrotation2010), or even deeper in time in the Mesozoic (lanWeakEquatorialSuperrotation2023), has further motivated study of this atmospheric phenomenon.
@@
Until now, the focus of superrotation research has been on the dynamical processes sustaining equatorial westerlies, as this requires non-trivial anti-diffusive angular momentum transport mechanisms (heldEquatorialSuperrotationEarthlike1999). Several routes have been suggested, all involving momentum transport by equatorial waves, either through enhanced diabatic heating (Kraucunas2005; arnoldAbruptTransitionStrong2012; Lutsko2018) or through a hydrodynamic instability (Williams2003; Mitchell2010; Potter2014; Zurita-Gotor2018a; Zurita-Gotor2022; Zurita-Gotor2024). It has further been suggested that those mechanisms might include positive feedbacks leading to an abrupt transition (arnoldAbruptTransitionStrong2012; herbertAtmosphericBistabilityAbrupt2020; Zurita-Gotor2022), which might make superrotation an example of a tipping point. Most of these dynamical studies were conducted in an idealized modeling framework of a dry atmosphere without explicitly taking into account processes including condensation of water vapor and radiation. Two exceptions are the studies of caballeroSpontaneousTransitionSuperrotation2010, which observed a spontaneous transition to superrotation under large radiative forcing, and arnoldEnhancedMJOlikeVariability2013, which reported weak superrotation in response to increasing sea surface temperature (SST). CaballeroCarlson2018 also investigated the plausibility of the scenarios suggested by pierrehumbertClimateChangeTropical2000 and tzipermanPlioceneEquatorialTemperature2009b by studying the possibility that tropical westerlies reached the surface.
@@
In this work, rather than focusing on the dynamical processes underlying superrotation, we instead explore the consequences that such a large-scale circulation change would have for climate if it were to occur, in particular for surface temperature and the water cycle. To do so, we impose superrotation using a prescribed torque in a moist model which includes parameterizations for major physical processes in the atmosphere such as radiation, convection, and condensation (see Section 2 for details). We find that superrotation is not restricted to a zonal wind change in the upper tropical troposphere but rather affects the general circulation as a whole, resulting in substantial impacts on surface temperature and precipitation. These impacts could potentially be important for interpreting paleoclimate proxies and the processes controlling warm climates in Earth’s past and future. The impact of (forced) superrotation on surface temperature, revealed in this study, together with the emergence of superrotation in simulations of hot climates  (caballeroSpontaneousTransitionSuperrotation2010; arnoldEnhancedMJOlikeVariability2013) suggests a possible feedback between superrotation and SST that has received little attention to date.
@@
After presenting the modeling setup used in our study (Section  2), we discuss the temperature and precipitation changes induced by superrotation (Section  3). We then build an understanding of the physical processes underlying the temperature changes by investigating the top-of-atmosphere (TOA) radiative budget using a forcing/feedback framework (Section  4) and reconstructing the simulated temperature response using an energy balance model (Section  5). Finally, we relate the superrotation-induced water cycle changes to the circulation changes (Section  6) before summarizing our results (Section  7).
@@ 2.1 Numerical experiments
In this study, we use the GCM Isca (vallisIscaV10Framework2018) in a relatively idealized moist aquaplanet configuration. The model, based on the GFDL dynamical core (Gordon1982), solves the moist primitive equations in vorticity-divergence form on 40 (unevenly spaced) $\sigma$ levels (Bourke1974) at T42 horizontal resolution.
@@
Latent heat of condensation and the radiative effects of water vapor are simulated, with the former represented through the large-scale condensation parameterization of Frierson2006. A standard Betts-Miller scheme for moist convection is also employed (bettsNewConvectiveAdjustment1986; bettsNewConvectiveAdjustment1986a). Radiative fluxes are computed using the comprehensive SOCRATES scheme (manners2017socrates), with a CO22 concentration of $300$ ppm and a modern ozone climatology. There is a diurnal cycle in the model but no seasonal cycle; a permanent equinox configuration is used for insolation. Clouds are not represented in our configuration of the model and the radiative effect of aerosols is not taken into account.
@@
The lower boundary is a global slab ocean with a $20$ m mixed layer depth. We impose prescribed meridional heat fluxes in the ocean, following Merlis2013. Sea ice is not represented in the model and the surface temperature is allowed to reach values below the freezing point. The surface albedo is prescribed to $0.25$ and standard drag laws govern the exchanges of momentum, sensible and latent heat at the surface, based on Monin-Obukhov theory as described in Frierson2006. At the top of the atmosphere (above $150~\rm{hPa}$), a sponge layer dissipates momentum with a Rayleigh drag of characteristic timescale $12~\rm{h}$.
@@
This type of idealized configuration (with variations regarding, for example, how radiative transfer is treated or whether prescribed or interactive SSTs are used) has been used extensively, not only as an evaluation framework for model intercomparison (Neale2000; Neale2000a; Lee2008) but also as a conceptual framework to connect theory with numerical models which remain interpretable (Blackburn2013). For instance, similar configurations have been used to study meridional energy transport (Caballero2005; Frierson2007c; Rose2012; Merlis2022) and the water cycle (OGorman2008; Byrne2016; Donohoe2019) over a range of climates, the response of the Hadley cells to various forcings (Levine2011; Merlis2013; hilgenbrinkResponseHadleyCirculation2018), convectively coupled waves (Frierson2007), and the annual cycle of the tropical tropopause layer (Jucker2017).
Here, we use the idealized configuration described above because it constitutes a relatively simple framework that nevertheless takes into account key feedbacks linking surface temperature and the water cycle to changes in large-scale circulation. The strategy we adopt to investigate the response of climate-relevant fields to the emergence of superrotation is to impose an atmospheric torque. We choose this numerical setup, rather than, for instance, letting superrotation emerge spontaneously from thermal forcings such as increased CO2 or sea surface temperatures, for two reasons. On the one hand, although the spontaneous emergence of superrotation has been reported for such thermal forcings in a small number of studies (caballeroSpontaneousTransitionSuperrotation2010; arnoldEnhancedMJOlikeVariability2013; lanWeakEquatorialSuperrotation2023), the transition occurs for very large forcing values and might be sensitive to parameterization choices (Zurita-Gotor2025). Hence it remains unclear to what extent it is robust and well represented by current models. In addition, our goal here is to isolate the climate changes, in particular the temperature and precipitation responses, resulting directly from the circulation changes associated with the transition to superrotation. This is more easily achieved with a forcing which does not directly affect these fields.
With this in mind, we impose a torque with a spatial structure corresponding to the convergence of the eddy zonal momentum fluxes generated by a zonally asymmetric diabatic heating, which has been used in a number of dynamical studies of superrotation (Kraucunas2005; arnoldAbruptTransitionStrong2012; Lutsko2018; herbertAtmosphericBistabilityAbrupt2020). This spatial structure of this imposed torque is the following:
$\mathbf{F}=F_{0}\frac{\left(1-\left(\frac{\phi}{\phi_{\mathrm{scale}}}\right)^{2}\right)}{\cos{\phi}}\exp\left(-\frac{1}{2}\left(\frac{\phi}{\phi_{\rm scale}}\right)^{2}\right)\exp\left(-\frac{1}{2}\left(\frac{p-p_{0}}{p_{\rm scale}}\right)^{2}\right)\mathbf{e}_{x},$ (1)
where $\phi_{\rm scale}=5^{\circ}$, $p_{0}=300$ hPa, $p_{\rm scale}=100$ hPa and $\mathbf{e}_{x}$ is the zonal unit vector. The torque has a tripolar structure, which is also similar to observed eddy momentum flux convergence associated to seasonal superrotation on Earth (Zhang2022). This structure integrates to $\pi\exp\left(-\frac{1}{2}\frac{\pi^{2}}{4\phi_{\rm scale}^{2}}\right)\approx 0$ on the meridional direction, which guarantees no net injection or damping of angular momentum. We vary the strength of the prescribed torque through the $F_{0}$ parameter, from 0 to 44 m.s1-1.day1-1. This parameter corresponds to the maximum value of the torque. While we use it to identify the different simulations, it should be kept in mind that typical values over regions of finite extent are smaller and similar to accelerations associated with the transition to superrotation reported elsewhere in the litterature. Note that eddy momentum flux convergence leading to superrotation could also be generated, for example, by a low-latitude mixed barotropic-baroclinic instability (Williams2003), or by a Kelvin-Rossby instability (Mitchell2010; Potter2014; Zurita-Gotor2018a; Zurita-Gotor2022; Zurita-Gotor2024), leading to eddy fluxes with a similar shape, albeit less focused on the tropics.
@@ 2.2 Forcing/feedback framework
We run several experiments with the model described above and interactive SSTs: a control run, with $F_{0}=0$ m.s1-1.day1-1, as well as forced runs with unit increments of $F_{0}$ up to $F_{0}=4$ m.s1-1.day1-1 (we will mostly focus on the $F_{0}=0$ and $F_{0}=4$ m.s1-1.day1-1 experiments). Analogous runs but with fixed SSTs are also performed: in those cases, we prescribe time-averaged SSTs from the control run. The fixed-SST experiments are run in order to diagnose the radiative forcing induced by the prescribed superrotation, in the standard way described for instance by Dessler2015. In particular, our aim is to decompose the changes in TOA radiation into radiative forcing and feedback components, which can each be broken down into contributions from different physical mechanisms. In this context, we can write the anomalous TOA energy budget as:
$\underbrace{\delta R}_{\text{Total TOA budget}}=\underbrace{F}_{\text{Radiative forcing}}-\underbrace{\lambda\delta T_{\rm s}}_{\text{Feedbacks}},$ (2)
@@
The radiative forcing term $F$ in Eq. (2), diagnosed using the fixed-SST simulations, can be interpreted as the change in the TOA radiative budget driven by the prescribed torque and resulting circulation change. The feedback term $\lambda\delta T_{\rm s}$, on the other hand, includes all changes in TOA radiation linearly related to surface temperature. This is similar to the standard framework used for studying climate sensitivity (sherwoodAssessmentEarthsClimate2020), with one notable difference. In the context of climate-sensitivity studies, the radiative forcing is driven only by the direct radiative effect of CO2 changes, with temperature and humidity held constant (in practice, estimating this radiative forcing precisely can be challenging due to ‘rapid adjustments’ of, for example, humidity to CO2 changes that are difficult to disentangle from the direct CO2 forcing (e.g., smith2018understanding)). In this study, the forcing does not directly affect the TOA radiative budget; instead the temperature and humidity profiles first adjust to the circulation changes at essentially constant SST, leading to a modified TOA radiative budget which we term the radiative forcing (this is analogous to the effective radiative forcing in the context of CO2 forcing (e.g., smith2020effective)). The slower adjustment of SST leads to further reorganization of the water vapor and temperature fields, which are included in the radiative feedback terms.
@@ 2.3 Radiative kernels
We decompose changes in the TOA radiation budget into components representing different physical processes using radiative kernels (sodenQuantifyingClimateFeedbacks2008). Radiative kernels represent the change of the radiative budget associated to a unit change of a given quantity at a given point in space. In general terms and assuming linearity, we can approximately write a TOA radiation change, $\delta R$, as a function of the radiative kernels and changes in the various physical quantities affecting TOA radiation: $\delta R\approx\sum_{X}K^{X}\delta X$, where $K^{X}=\partial R/\partial X$ is a radiative kernel and $X$ represents quantities affecting the TOA radiation. In the configuration considered here, there are three relevant radiative kernels, $K^{T_{\mathrm{s}}}=\left.\frac{\partial R}{\partial T_{\mathrm{s}}}\right|_{T,q}$, $K^{T}=\left.\frac{\partial R}{\partial T}\right|_{T_{\mathrm{s}},q}$ and $K^{\mathrm{wv}}=\left.\frac{\partial R}{\partial\log(q)}\right|_{T_{\mathrm{s}},T}\frac{\partial\log(q)}{\partial T}$, quantifying the sensitivity of the TOA radiative budget to changes in surface temperature, atmospheric temperature and specific humidity, respectively (there are no changes in chemical species concentrations or albedo in our simulations).
In the analysis below, we will use these kernels to decompose the changes of TOA radiative budget, but we will use relative humidity as a variable instead of specific humidity, so that the total TOA radiative budget changes write $\delta R_{\mathrm{tot,kernel}}=\delta R_{T_{\mathrm{s}}}+\delta R_{\Gamma}+\delta R_{\mathrm{RH}}$, with $\delta R_{T_{\mathrm{s}}}=(K^{T_{\mathrm{s}}}+K^{T}+K^{\mathrm{wv}})\delta T_{\mathrm{s}}$, $\delta R_{\Gamma}=(K^{T}+K^{\mathrm{wv}})(\delta T-\delta T_{\mathrm{s}})$ and $\delta R_{\mathrm{RH}}=K^{\mathrm{wv}}\frac{\delta\log\left(\mathrm{RH}\right)}{\partial\log(q)/\partial T}$ (see Appendix B for a derivation).
In practice, radiative kernels are computed by running a radiation scheme on two base states that differ only by a local change in the chosen variable, for instance a $1$ K temperature change or a $1$ % specific humidity change at a given point in space. Here, we use kernels previously computed for Isca using the same setup as that employed in this study (aquaplanet, no clouds or sea ice, T4242 resolution, $40$ vertical levels) (liuRadiativeKernelsIsca2020), but with the RRTM radiative scheme (mlawerRadiativeTransferInhomogeneous1997) (rather than SOCRATES). To validate this approach, we have verified that the total radiative budget computed from the kernels approximately matches the simulated radiative budget from the experiments (see Section 4.4.2).
Note that, as mentioned above, unlike the more familiar case of CO2 forcing, here both the radiative forcing and the radiative feedback terms introduced in Section 22.2 can be decomposed using the radiative kernels: we can write $\delta R=\sum_{X}\delta R_{X}=F-\lambda\delta T_{s}=\sum_{X}F_{X}-(\lambda\delta T_{S})_{X}$, where the sum runs over the surface temperature, air temperature and humidity fields.